The Alford Group had 240,000 shares of common stock outstanding at January 1, 2018. The following activities affected common shares during the year. There are no potential common shares outstanding.
  

2018
Feb. 28 Purchased 30,000 shares of treasury stock.
Oct. 31 Sold the treasury shares purchased on February 28.
Nov. 30 Issued 120,000 new shares.
Dec. 31 Net income for 2018 is $1,495,000.
2019
Jan. 15 Declared and issued a 2-for-1 stock split.
Dec. 31 Net income for 2019 is $1,495,000.


Required:

1. Determine the 2018 EPS.
2. Determine the 2019 EPS.
3. At what amount will the 2018 EPS be presented in the 2019 comparative financial statements?
Answer #1
1. EPS 2018 $6.5 per share  
2. EPS 2019 $2.08 per share  
3. EPS 2018 to be presented in 2019 financial statement $3.25 per share  

Working:

1. EPS (2018)
Weighted average number of shares :
Weighted average number of shares =(240000-(30000*10/12)+(30000*2/12)+(120000*1/12))shares  
Weighted average number of shares 230,000 shares  
Net Income                                       1,495,000
EPS = Net Income / Weighted Average Outstanding Shares =$1495000/230000 shares = $6.5 per share
2. EPS (2019)
Weighted average number of shares = number of shares outstanding on Jan 15, 2019 * 2 = (240000-30000+30000+120000) * 2   
Weighted average number of shares 720,000 shares  
Net Income                                       1,495,000
EPS = Net Income / Weighted Average Outstanding Shares =$1495000/720000 shares = $2.08 per share
3. 2018 EPS to be presented in the 2019 comparative financial statements:
Weighted average number of shares of 2018 now adjusted for stock split = 230000 x 2 = 460000.
EPS = $ 1,495000 / 460000 = $ 3.25

Please give positive rating.